OpenStudy (anonymous):

which equations have at least one positive solution for x: -x^2+6x+2=0 -x^2+6x-8=0 1/4x^2+3x+8=0 1/4x^2+3x+12=0

OpenStudy (anonymous):

-x^2+6x+2=0 (x = 3+sqrt(11) ~~ 6.31662) -x^2+6x-8=0 (x=4 ,x=2)